OpenStudy (anonymous):

4x+4/2 = 10 what is X?

OpenStudy (anonymous):

sorry i didnt make that clear first time

OpenStudy (vjohnny):

x=2, [ 4x+2=10 ==> 4x=10-2 ==> 4x=8 ==> x=8/4 ==> x=2]

OpenStudy (anonymous):

thanks!